# SABRE-FL: Selective and Accurate Backdoor Rejection for Federated Prompt Learning

### v1

@@ 1 Introduction
Federated Learning (FL) mcmahan2017communication enables decentralized model training across multiple users while keeping data local, thereby preserving privacy and reducing centralized risks. In FL, clients independently train models on local data and share only model updates with a server, which aggregates them into a global model. Due to its privacy-preserving nature, FL has been adopted in settings like Google’s Gboard gboard for next-word prediction, Apple’s Siri technologyreviewApplePersonalizes for automatic speech recognition, and WeBank for credit risk prediction webankcredit . Recent advances have extended FL to support more expressive models, such as vision-language models, by integrating prompt-based learning zhou2022learning ; guo2023promptfl ; deng2024unlocking .
Prompt learning is a recent paradigm that adapts large pre-trained models such as OpenAI’s CLIP (Contrastive Language-Image Pretraining) radford2021learning to downstream tasks by optimizing lightweight, learnable input prompts instead of finetuning the full model. Originally developed in centralized settings, prompt learning has shown impressive few-shot generalization, task transferability, and reduced compute cost, particularly with vision-language models zhou2022learning ; zhou2022conditional . Motivated by these advantages, recent works have introduced prompt learning into FL guo2023promptfl ; weng2024probabilistic , giving rise to federated prompt learning (FPL). In FPL, clients independently optimize prompt vectors while keeping the model backbone frozen, and share only these prompts with the server. This design greatly reduces communication and memory overhead and enables efficient cross-client adaptation in multimodal and heterogeneous environments.
Despite its appeal, FL is not inherently secure kairouz2019advances . In practice, some clients may behave maliciously, either by corrupting their local training data or manipulating model updates, to influence the behavior of the global model fang2020local ; shejwalkar2022back ; wang2020attack ; shejwalkar2021manipulating ; baruch2019a ; mhamdi2018the ; xie2019fall ; bhagoji2019analyzing ; bagdasaryan2018how . A particularly insidious example is the backdoor attack nguyen2023iba ; zhang2023a3fl ; bagdasaryan2020backdoor ; wang2020attack , in which an adversary injects carefully crafted inputs (called triggers) into local training data (Figure 1). These triggers cause the global model to misclassify specific test-time inputs while preserving high accuracy on benign samples. Prior work on backdoor attacks in FL has largely focused on traditional classification tasks in unimodal settings nguyen2023iba ; zhang2023a3fl , leaving the security properties of multimodal and prompt-based FL systems underexplored.
@@
While prompt learning in FL is gaining momentum, its security properties remain largely unexamined. This raises a key question: how vulnerable is Federated Prompt Learning to backdoor attacks? In this work, we show that prompt learners in FL are highly susceptible to backdoors, even when model updates are limited to prompt vectors. We introduce a backdoor attack that inserts a learnable, visually imperceptible trigger into a subset of clients’ training data. The attack draws inspiration from BadClip bai2024badclip who design a trigger-based backdoor attack for prompt-learning in the centralized setting. In our attack each malicious client has its own malicious trigger that pushes the prompt embeddings toward a target label in CLIP’s semantic space, leading to high-confidence misclassification at inference. The attack remains stealthy and retains high clean accuracy across clients, matching performance observed in centralized prompt tuning. This demonstrates that Federated Prompt Learning is vulnerable to trigger-based backdoor attacks even when a few clients act maliciously. To the best of our knowledge, we are the first to study backdoor attacks in this setting, i.e., trigger-based attacks in multimodal federated prompt learning.
@@ 2.1 Federated Learning (FL)
In FL kairouz2019advances ; mcmahan2017communication , a central entity, known as the server, aims to train a global model, $\theta^{g}$, using private data distributed across multiple clients, without directly accessing their data. In each communication round, the server selects $n$ out of $N$ available clients and sends them the current global model $\theta^{t}_{g}$, where $t$ denotes the round index. Each selected client $k$ computes an update $\nabla^{t}_{k}$ using its local dataset $D_{k}$, and returns it to the server, which aggregates all updates using a predefined aggregation rule, such as FedAvg mcmahan2017communication .
In FedAvg, a client $k$ fine-tunes $\theta^{t}_{g}$ on their local data using stochastic gradient descent (SGD) for a fixed number of local epochs $E$, resulting in an updated local model $\theta^{t}_{k}$. The client then computes their update as the difference $\nabla^{t}_{k}=\theta^{t}_{k}-\theta^{t}_{g}$ and shares $\nabla^{t}_{k}$ with the server. Next, the server computes an aggregate of client updates, $f_{\mathsf{agg}}$ using mean, i.e.,
$\nabla^{t}_{\mathsf{agg}}=f_{\mathsf{mean}}(\nabla^{t}_{\{k\in[n]\}}).$ (1)
The server then updates the global model of the $(t+1)^{th}$ round using SGD and server learning $\eta$ as:
$\theta^{t+1}_{g}\leftarrow\theta^{t}_{g}+\eta\nabla^{t}_{\mathsf{agg}}$ (2)
@@ 2.2 Prompt Learning with Vision-Language Models
Vision-Language Models: Large vision-language models (VLMs), such as CLIP radford2021learning , have demonstrated remarkable generalization across diverse downstream tasks. By aligning images and text in a shared semantic space, these models enable strong zero-shot and few-shot performance without task-specific supervision. However, their size, often exceeding hundreds of millions of parameters, makes traditional fine-tuning computationally expensive and bandwidth-intensive, particularly in distributed or resource-constrained environments.
Prompt Learning zhou2022learning : Prompt learning adapts large pre-trained models to downstream tasks by introducing a set of learnable prompt vectors that are prepended to the model input. During training, only these prompts are updated, allowing efficient task adaptation while keeping the backbone frozen. This reduces the number of trainable parameters and computational cost, making the approach particularly attractive for few-shot and resource-constrained settings. Prompt learning has been shown to be effective across multiple modalities khattak2022maple ; zhou2022conditional ; zhou2022learning . In CLIP-based architectures, this involves optimizing a set of context vectors $\bm{V}=[\bm{v}_{1},\bm{v}_{2},\dots,\bm{v}_{N}]^{\top}\in\mathbb{R}^{N\times e}$, where each $\bm{v}_{i}$ is a learnable token embedding and $e$ is the embedding dimension. Given an input image $\bm{x}$ and a class name embedding $\bm{c}_{i}$, the image encoder $f(\bm{x})$ and the text encoder $g(\{\bm{V},\bm{c}_{i}\})$ produce modality-aligned representations. The prediction probability is computed using cosine similarity:
@@
Prompt Learning in FL: The benefits of prompt learning mentioned above have motivated its integration into the federated setting guo2023promptfl ; guo2023pfedprompt ; zhao2023fedprompt . In federated prompt learning, each client optimizes a local prompt vector while keeping the foundation model, e.g., CLIP, frozen, and transmits only the prompt to the server for aggregation. This substantially reduces memory usage and communication cost, making it feasible to deploy foundation models like CLIP in privacy-preserving, bandwidth-limited environments. Such systems have demonstrated strong downstream performance across vision and multimodal tasks while maintaining FL’s privacy and scalability benefits.
@@ 2.3 Backdoor Attacks
Backdoor attacks  bai2021targeted ; bai2022hardly ; bai2023versatile ; gao2023imperceptible ; li2021invisible ; turner2019label ; xu2024towards ; ya2024towards are a class of training-time data poisoning techniques wherein an adversary injects carefully crafted samples into the training set with the goal of inducing targeted misbehavior at test time bagdasaryan2018how ; hanif2024baple . These poisoned samples contain an imperceptible or benign-looking trigger, such as a small patch in the input, and are assigned a target label of the attacker’s choosing. Once trained on such data, the compromised model behaves normally on clean inputs but misclassifies any input containing the trigger to the target class. This selective misbehavior makes backdoor attacks particularly insidious, as they are challenging to detect using standard validation procedures. Backdoor attacks have been studied across multiple modalities including vision gu2017badnets , language kurita2020weight , and multimodal models liang2024badclip ; bai2024badclip and have proven effective even in privacy-preserving settings like FL, where model updates rather than raw data are shared.
Backdoor Attacks in Federated Learning: Backdoor attacks pose a serious security threat to FL, allowing adversaries to embed malicious behavior into the global model by manipulating a small number of clients during training bagdasaryan2020backdoor ; wang2020attack ; shejwalkar2021manipulating . These attacks typically preserve high accuracy on clean inputs while causing targeted misclassification on inputs containing an attacker-defined trigger. Early approaches rely on fixed triggers sun2019can ; xie2019dba ; bhagoji2019analyzing , while more recent methods optimize trigger patterns to maximize attack success and stealthiness nguyen2023iba ; zhang2023a3fl . For example, A3FL zhang2023a3fl predicts the movement of the global model updates and improves attack durability by ensuring the backdoor persists across global aggregation rounds. Similarly, IBA nguyen2023iba jointly optimizes a visually stealthy trigger and selectively poisons models’ parameters that are less likely to be updated by the main task’s learning process, achieving a durable and stealthy backdoor effect.
@@ 3.1 Overview
While backdoor attacks have been extensively studied in traditional unimodal FL settings, their feasibility in multimodal federated prompt learning remains underexplored. Unlike traditional full-model FL, it exposes a narrower attack surface, limited to lightweight prompt vectors, raising new questions about the strength, persistence, and stealth of such attacks. These differences motivate our central hypothesis.
@@
Theoretical Motivation. In CLIP-based prompt learning zhou2022learning ; zhou2022conditional , classification is based on the cosine similarity between an image embedding $f(\bm{x})$ and a prompt-conditioned text embedding $g(\{\bm{V},\bm{c}_{i}\})$ for class $i$. To induce targeted misclassification toward a specific class $t$, it suffices to craft an input $\bm{x}^{\star}$ such that:
@@
This condition ensures that the model classifies $\bm{x}^{\star}$ as belonging to the target class $t$. In practice, our attack injects a visually imperceptible trigger, as shown in Figure 1, into local training data and optimizes it to shift image embeddings toward $g(\{\bm{V},\bm{c}_{t}\})$, effectively planting a backdoor in the global prompt learner. While this idea is inspired by prior work on backdoor optimization bai2024badclip ; zhang2023a3fl ; nguyen2023iba , adapting it to the prompt-only FL setting introduces new challenges: the global model is now updated solely via lightweight prompt vectors, and the image encoder remains frozen. This means the backdoor signal must propagate indirectly through prompt aggregation, requiring the trigger to consistently bias prompt updates without direct influence over model weights, making the optimization problem both weaker in signal and more sensitive to noise. The attack is visually explained in Figure 1.
@@ 3.2 Threat Model
Capabilities. We assume a standard FL setup with $N$ clients and a central server aggregating client prompt updates. The adversary controls a fraction $m/N$ of clients, set to 25% by default, consistent with prior works cao2022fedrecover ; cao2020fltrust . The attacker can:
Knowledge. Since the attacker controls client devices, it naturally has access to the full prompt learning setup, including model architecture, frozen CLIP backbone, and training procedure. This is a standard assumption in federated backdoor attack literature bagdasaryan2020backdoor ; shejwalkar2022back , and reflects realistic adversaries in open-source or distributed deployments where models like CLIP are publicly available radford2021learning .
@@ 3.3 Design of the Backdoor Attack in an FL Setting
Table 1: Backdoor attack effectiveness without (no attack) and with (clean & backdoor) attack.
@@ 3.4 Attack Impact
Comparison with Centralized Backdoor Attacks: We compare our FL backdoor attack against its centralized counterpart, BadCLIP bai2024badclip , which serves as the baseline for prompt-learning backdoor attacks in non-federated settings. BadCLIP achieves near 100% backdoor success by directly poisoning a large portion of the training data and optimizing the trigger in a fully centralized regime. In contrast, our setting uses the standard FedAvg aggregation algorithm and models a more realistic adversary: only a small subset of clients are malicious, and poisoning is confined to local updates. This naturally dilutes the backdoor signal during aggregation and results in lower backdoor accuracy compared to the centralized case. Despite this, our attack achieves high success rates on several datasets, demonstrating that prompt-based FL remains vulnerable even with limited adversarial participation. In Table 1, we report results under the no-defense scenario to highlight how much damage can occur with the default FedAvg setup. We analyze the effectiveness of standard defenses in mitigating this attack later in §5.2.

### v2

@@ 1 Introduction
Federated Learning (FL) McMahan et al. (2017) enables decentralized model training across multiple users while keeping data local, thereby preserving privacy and reducing centralized risks. In FL, clients independently train models on local data and share only model updates with a server, which aggregates them into a global model. Due to its privacy-preserving nature, FL has been adopted in settings like Google’s Gboard 22 for next-word prediction, Apple’s Siri 31 for automatic speech recognition, and WeBank for credit risk prediction webankcredit (2019). Recent advances have extended FL to support more expressive models, such as vision-language models, by integrating prompt-based learning Zhou et al. (2022b); Guo et al. (2023a); Deng et al. (2024).
Prompt learning is a recent paradigm that adapts large pre-trained models such as OpenAI’s CLIP (Contrastive Language-Image Pretraining) Radford et al. (2021) to downstream tasks by optimizing lightweight, learnable input prompts instead of finetuning the full model. Originally developed in centralized settings, prompt learning has shown impressive few-shot generalization, task transferability, and reduced compute cost, particularly with vision-language models Zhou et al. (2022b; a). Motivated by these advantages, recent works have introduced prompt learning into FL Guo et al. (2023a); Weng et al. (2024), giving rise to federated prompt learning (FPL). In FPL, clients independently optimize prompt vectors while keeping the model backbone frozen, and share only these prompts with the server. This design greatly reduces communication and memory overhead and enables efficient cross-client adaptation in multimodal and heterogeneous environments.
Despite its appeal, FL is not inherently secure Kairouz et al. (2019). In practice, some clients may behave maliciously, either by corrupting their local training data or manipulating model updates, to influence the behavior of the global model Fang et al. (2020); Shejwalkar et al. (2022); Wang et al. (2020); Shejwalkar and Houmansadr (2021); Baruch et al. (2019); Mhamdi et al. (2018); Xie et al. (2019b); Bhagoji et al. (2019); Bagdasaryan et al. (2020a). A particularly insidious example is the backdoor attack Nguyen et al. (2023); Zhang et al. (2023); Bagdasaryan et al. (2020b); Wang et al. (2020), in which an adversary injects carefully crafted inputs (called triggers) into local training data (Figure 1). These triggers cause the global model to misclassify specific test-time inputs while preserving high accuracy on benign samples. Prior work on backdoor attacks in FL has largely focused on traditional classification tasks in unimodal settings Nguyen et al. (2023); Zhang et al. (2023), leaving the security properties of multimodal and prompt-based FL systems underexplored.
@@
While prompt learning in FL is gaining momentum, its security properties remain largely unexamined. This raises a key question: how vulnerable is Federated Prompt Learning to backdoor attacks? In this work, we show that prompt learners in FL are highly susceptible to backdoors, even when model updates are limited to prompt vectors. We introduce a backdoor attack that inserts a learnable, visually imperceptible trigger into a subset of clients’ training data. The attack draws inspiration from BadClip Bai et al. (2024) who design a trigger-based backdoor attack for prompt-learning in the centralized setting. In our attack each malicious client has its own malicious trigger that pushes the prompt embeddings toward a target label in CLIP’s semantic space, leading to high-confidence misclassification at inference. The attack remains stealthy and retains high clean accuracy across clients, matching performance observed in centralized prompt tuning. This demonstrates that Federated Prompt Learning is vulnerable to trigger-based backdoor attacks even when a few clients act maliciously. To the best of our knowledge, we are the first to study backdoor attacks in this setting, i.e., trigger-based attacks in multimodal federated prompt learning.
@@ 2.1 Federated Learning (FL)
In FL Kairouz et al. (2019); McMahan et al. (2017), a central entity, known as the server, aims to train a global model, $\theta^{g}$, using private data distributed across multiple clients, without directly accessing their data. In each communication round, the server selects $n$ out of $N$ available clients and sends them the current global model $\theta^{t}_{g}$, where $t$ denotes the round index. Each selected client $k$ computes an update $\nabla^{t}_{k}$ using its local dataset $D_{k}$, and returns it to the server, which aggregates all updates using a predefined aggregation rule, such as FedAvg McMahan et al. (2017).
In FedAvg, a client $k$ fine-tunes $\theta^{t}_{g}$ on their local data using stochastic gradient descent (SGD) for a fixed number of local epochs $E$, resulting in an updated local model $\theta^{t}_{k}$. The client then computes their update as the difference $\nabla^{t}_{k}=\theta^{t}_{k}-\theta^{t}_{g}$ and shares $\nabla^{t}_{k}$ with the server. Next, the server computes an aggregate of client updates, $f_{\mathsf{agg}}$ using mean, i.e., $\nabla^{t}_{\mathsf{agg}}=f_{\mathsf{mean}}(\nabla^{t}_{\{k\in[n]\}})$ and updates the global model of the $(t+1)^{th}$ round using SGD and server learning $\eta$ as: $\theta^{t+1}_{g}\leftarrow\theta^{t}_{g}+\eta\nabla^{t}_{\mathsf{agg}}$.
@@ 2.2 Prompt Learning with Vision-Language Models
Vision-Language Models: Large vision-language models (VLMs), such as CLIP Radford et al. (2021), have demonstrated remarkable generalization across diverse downstream tasks. By aligning images and text in a shared semantic space, these models enable strong zero-shot and few-shot performance without task-specific supervision. However, their size, often exceeding hundreds of millions of parameters, makes traditional fine-tuning computationally expensive and bandwidth-intensive, particularly in distributed or resource-constrained environments.
Prompt Learning Zhou et al. (2022b): Prompt learning adapts large pre-trained models to downstream tasks by introducing a set of learnable prompt vectors that are prepended to the model input. During training, only these prompts are updated, allowing efficient task adaptation while keeping the backbone frozen. This reduces the number of trainable parameters and computational cost, making the approach particularly attractive for few-shot and resource-constrained settings. Prompt learning has been shown to be effective across multiple modalities Khattak et al. (2022); Zhou et al. (2022a; b). In CLIP-based architectures, this involves optimizing a set of context vectors $\bm{V}=[\bm{v}_{1},\bm{v}_{2},\dots,\bm{v}_{N}]^{\top}\in\mathbb{R}^{N\times e}$, where each $\bm{v}_{i}$ is a learnable token embedding and $e$ is the embedding dimension. Given an input image $\bm{x}$ and a class name embedding $\bm{c}_{i}$, the image encoder $f(\bm{x})$ and the text encoder $g(\{\bm{V},\bm{c}_{i}\})$ produce modality-aligned representations. The prediction probability is computed using cosine similarity:
@@
Prompt Learning in FL: The benefits of prompt learning mentioned above have motivated its integration into the federated setting Guo et al. (2023a; b); Zhao et al. (2023). In federated prompt learning, each client optimizes a local prompt vector while keeping the foundation model, e.g., CLIP, frozen, and transmits only the prompt to the server for aggregation. This substantially reduces memory usage and communication cost, making it feasible to deploy foundation models like CLIP in privacy-preserving, bandwidth-limited environments. Such systems have demonstrated strong downstream performance across vision and multimodal tasks while maintaining FL’s privacy and scalability benefits.
@@ 2.3 Backdoor Attacks
Backdoor attacks  Bai et al. (2021; 2022; 2023a); Gao et al. (2023); Li et al. (2021); Turner et al. (2019); Xu et al. (2024); Ya et al. (2024) are a class of training-time data poisoning techniques wherein an adversary injects carefully crafted samples into the training set with the goal of inducing targeted misbehavior at test time Bagdasaryan et al. (2020a); Hanif et al. (2024). These poisoned samples contain an imperceptible or benign-looking trigger, such as a small patch in the input, and are assigned a target label of the attacker’s choosing. Once trained on such data, the compromised model behaves normally on clean inputs but misclassifies any input containing the trigger to the target class. This selective misbehavior makes backdoor attacks particularly insidious, as they are challenging to detect using standard validation procedures. Backdoor attacks have been studied across multiple modalities including vision Gu et al. (2017), language Kurita et al. (2020), and multimodal models Liang et al. (2024); Bai et al. (2024) and have proven effective even in privacy-preserving settings like FL, where model updates rather than raw data are shared.
Backdoor Attacks in Federated Learning: Backdoor attacks pose a serious security threat to FL, allowing adversaries to embed malicious behavior into the global model by manipulating a small number of clients during training Bagdasaryan et al. (2020b); Wang et al. (2020); Shejwalkar and Houmansadr (2021). These attacks typically preserve high accuracy on clean inputs while causing targeted misclassification on inputs containing an attacker-defined trigger. Early approaches rely on fixed triggers Sun et al. (2019); Xie et al. (2019a); Bhagoji et al. (2019), while more recent methods optimize trigger patterns to maximize attack success and stealthiness Nguyen et al. (2023); Zhang et al. (2023). For example, A3FL Zhang et al. (2023) predicts the movement of the global model updates and improves attack durability by ensuring the backdoor persists across global aggregation rounds. Similarly, IBA Nguyen et al. (2023) jointly optimizes a visually stealthy trigger and selectively poisons models’ parameters that are less likely to be updated by the main task’s learning process, achieving a durable and stealthy backdoor effect.
@@ 3.1 Overview
While backdoor attacks have been extensively studied in traditional unimodal FL settings, their feasibility in multimodal federated prompt learning remains underexplored. Unlike traditional full-model FL, it exposes a narrower attack surface, limited to prompt vectors, raising new questions about strength, persistence, and stealth of such attacks. These differences motivate our central hypothesis.
@@
Positioning our work relative to existing literature. Several recent works have demonstrated the vulnerability of contrastive vision-language models like CLIP to backdoor attacks in centralized settings. Notably, BadCLIP Bai et al. (2024) introduces a powerful trigger-aware attack that jointly manipulates both the image and text encoders using prompt-conditioned triggers. A similar variant Liang et al. (2024) improves stealth and robustness using dual-embedding alignment. Other works such as BadEncoder Jia et al. (2022) and contrastive poisoning attacks Carlini and Terzis (2021) inject backdoors directly into frozen image encoders or pretraining datasets. While effective, these attacks are designed for centralized or pretraining regimes. In this paper, we focus exclusively on the BadCLIP attack due to its compatibility with prompt tuning and its relevance to the downstream FL scenario explored in our work.
Theoretical Motivation. In CLIP-based prompt learning Zhou et al. (2022b; a), classification is based on the cosine similarity between an image embedding $f(\bm{x})$ and a prompt-conditioned text embedding $g(\{\bm{V},\bm{c}_{i}\})$ for class $i$. To induce targeted misclassification toward a specific class $t$, it suffices to craft an input $\bm{x}^{\star}$ such that:
@@
This condition ensures that the model classifies $\bm{x}^{\star}$ as belonging to the target class $t$. In practice, our attack injects a visually imperceptible trigger, as shown in Figure 1, into local training data and optimizes it to shift image embeddings toward $g(\{\bm{V},\bm{c}_{t}\})$, effectively planting a backdoor in the global prompt learner. While this idea is inspired by prior work on backdoor optimization Bai et al. (2024); Zhang et al. (2023); Nguyen et al. (2023), adapting it to the prompt-only FL setting introduces new challenges: the global model is now updated solely via lightweight prompt vectors, and the image encoder remains frozen. This means the backdoor signal must propagate indirectly through prompt aggregation, requiring the trigger to consistently bias prompt updates without direct influence over model weights, making the optimization problem both weaker in signal and more sensitive to noise. The attack is visually explained in Figure 1.
@@ 3.2 Threat Model
Capabilities. We assume a standard FL setup with $N$ clients and a central server aggregating client prompt updates. The adversary controls a fraction $m/N$ of clients, set to 25% by default, consistent with prior works Cao et al. (2023; 2021). The attacker can:
  • Modify a subset of local training data by adding a learnable backdoor trigger to inputs.
  • Relabel triggered samples to the desired target class, known in literature as dirty-label attack Shejwalkar et al. (2023); Gu et al. (2019); Chen et al. (2017); Sarkar et al. (2021); Zeng et al. (2021); Nguyen and Tran (2020); Li et al. (2021).
  • Optimize the trigger jointly with the prompt learner at each malicious client to maximize its effect on the global prompt vector.
Knowledge. Since the attacker controls client devices, it naturally has access to the full prompt learning setup, including model architecture, frozen CLIP backbone, and training procedure. This is a standard assumption in federated backdoor attack literature Bagdasaryan et al. (2020b); Shejwalkar et al. (2022), and reflects realistic adversaries in open-source or distributed deployments where models like CLIP are publicly available Radford et al. (2021).
@@ 3.3 Design of the Backdoor Attack in an FL Setting
Table 1: Accuracy with no attack (clean model), clean inputs under attack, and backdoored inputs.
@@ 3.4 Attack Impact
Comparison with Centralized Backdoor Attacks: We compare our FL backdoor attack against its centralized counterpart, BadCLIP Bai et al. (2024), which serves as the baseline for prompt-learning backdoor attacks in non-federated settings. BadCLIP achieves near 100% backdoor success by directly poisoning a large portion of the training data and optimizing the trigger in a fully centralized regime. In contrast, our setting uses the standard FedAvg aggregation algorithm and models a more realistic adversary: only a small subset of clients are malicious, and poisoning is confined to local updates. This naturally dilutes the backdoor signal during aggregation and results in lower backdoor accuracy compared to the centralized case. Despite this, our attack achieves high success rates on several datasets, demonstrating that prompt-based FL remains vulnerable even with limited adversarial participation. In Table 1, we report results under the no-defense scenario to highlight how much damage can occur with the default FedAvg setup. We analyze the effectiveness of standard defenses in mitigating this attack later in §5.1.